A man invested $6,000, part of it at 5% simple interest and the rest at 7% simple interest. If his annual interest income is $372, how much did he invest at each rate?
Answer by Earlsdon(6287)

You can set up your equation like this: Let x = the amount invested at 5% simple interest.
5%(x) + 7%($6,000.00-x) = $372.00 Simplify and solve for x. Change percents to decimals.
0.05(x) + 0.07($6,000.00-x) = $372.00
0.05x + $420.00 - 0.07x = $372.00 Subtract $420.00 from both sides.
(0.05-0.07)x = -$48.00
-0.02x = -$48.00 Divide both sides by -0.02
x = $2,400.00 invested at 5% simple interest.
$6,000.00-$2,400.00 = $3,600.00 invested at 7% simple interest.
Check:
0.05($2,400.00) + 0.07($3,600) = ? If it comes to $372.00, all is cool!